OpenStudy (anonymous):

8x-9=3x+16 find value of x. What do I do first again?

OpenStudy (anonymous):

multiply

OpenStudy (paki):

add similar terms.. .then solve it by BODMAS rule...

OpenStudy (nurali):

8x-9=3x+16 Add both side by 9 8x-9+9=3x+16+9 8x=3x+25 Subtract both side by 3x 8x-3x=3x-3x+25 5x=25 Divide both side by 5 5x/5=25/5 x=5

OpenStudy (paki):

*combine